# Temporal Test-Time Adaptation with State-Space Models

### v2

@@ Abstract
Distribution shifts between training and test data are inevitable over the lifecycle of a deployed model, leading to performance decay. Adapting a model on test samples can help mitigate this drop in performance. However, most test-time adaptation methods have focused on synthetic corruption shifts, leaving a variety of distribution shifts underexplored. In this paper, we focus on distribution shifts that evolve gradually over time, which are common in the wild but challenging for existing methods, as we show. To address this, we propose STAD, a probabilistic state-space model that adapts a deployed model to temporal distribution shifts by learning the time-varying dynamics in the last set of hidden features. Without requiring labels, our model infers time-evolving class prototypes that act as a dynamic classification head. Through experiments on real-world temporal distribution shifts, we show that our method excels in handling small batch sizes and label shift.
@@ 1 Introduction
Synthetically corrupted images (e.g. CIFAR-10-C) are by far the most commonly used benchmark for assessing progress on TTA—with previous work noting a lack of benchmark diversity (Zhao et al., 2023b). These shifts increase the degree of information loss over time, and well-performing TTA methods must learn to preserve a static underlying signal. In this work we focus on a distribution shift of quite a different nature: Temporal distribution shifts are due to factors endogenous to the environment and often encode structural change, not just information loss. For example, we will experiment with the functional map of the world (FMoW) benchmark, which has the goal of classifying how land is used as it is developed over time (e.g. rural to urban). Gradual structural change over time has been studied in related subfields, such as temporal domain generalization (Bai et al., 2023), yet received little attention in TTA. We show that this setting, temporal test-time adaptation (TempTTA), poses significant challenges for existing TTA methods.
To address this gap, we propose State-space Test-time Adaptation (STAD), a method that builds on the power of probabilistic state-space models (SSMs) to represent non-stationary data distributions over time. STAD dynamically adapts a model’s final layer to accommodate an evolving test distribution. Specifically, we employ a SSM to track the evolution of the weight vectors in the final layer, where each vector represents a class, as distribution shift occurs. For generating predictions on newly acquired test batches, we use the SSM’s updated cluster means as the new parameters. STAD leverages Bayesian updating and does not rely upon normalization mechanisms. As a consequence, STAD excels in scenarios where many TTA methods collapse (Niu et al., 2023), such as adapting with very few samples and handling online class imbalance. Our contributions are the following:
  • In Sec. 2, we detail the setting of temporal test-time adaptation (TempTTA), which aims to cope with shifts that gradually evolve due to variation in the application domain. Despite being ubiquitous in real-world scenarios, these shifts are understudied in the TTA literature and pose significant challenges to established methods, as we demonstrate in Sec. 5.1 (Tab. 2).
  • In Sec. 3, we propose STAD, a novel method for TempTTA. It adapts to temporal distribution shifts by modeling its dynamics in representation space. No previous work has explicitly modeled these dynamics, which we demonstrate is crucial via an ablation study (Tab. 5).
  • In Sec. 5, we conduct a comprehensive evaluation of STAD and prominent TTA baselines under authentic temporal shifts. Our results show that STAD provides persistent performance gains, particularly in the cases of label shift and small batch sizes.
@@ 2 Problem Setting
We focus on the traditional setting of multi-class classification, where $\mathcal{X}\subseteq{\mathbb{R}}^{D}$ denotes the input (feature) space and $\mathcal{Y}\subseteq\{1,\dots,K\}$ denotes the label space. Let ${\mathbf{x}}$ and y be random variables and $\mathbb{P}({\mathbf{x}},{\textnormal{y}})=\mathbb{P}({\mathbf{x}})\ \mathbb{P}({\textnormal{y}}|{\mathbf{x}})$ the unknown source data distribution. We assume ${\bm{x}}\in\mathcal{X}$ and $y\in\mathcal{Y}$ are realisations of ${\mathbf{x}}$ and y. The goal of classification is to find a mapping $f_{\theta}$, with parameters ${\theta}$, from the input space to the label space $f_{\theta}:\mathcal{X}\to\mathcal{Y}$. Fitting the classifier $f_{\theta}$ is usually accomplished by minimizing an appropriate loss function (e.g. log loss). Yet, our method is agnostic to how $f_{\theta}$ is trained and therefore easy to use with, for instance, a pre-trained model from the web.
Table 1: TempTTA compared
source data free evolving test dist. time index bounded shift
TDG
TTA
CTTA
TempTTA
@@
We are interested in adapting a model at test-time to a test distribution that evolves with time. Such temporal distribution shifts have been the focus of work in temporal domain generalization (TDG) (Nasery et al., 2021; Qin et al., 2022; Bai et al., 2023; Zeng et al., 2024). Our setting differs from TDG in that we do not require access to the source data and do not alternate the training procedure. While previous work on continual test-time adaptation (CTTA) study adaptation to changing target domains, they mostly discrete domains (e.g. different corruption types). e consider a special case of CTTA that is defined by two key aspects: the domain index is temporal and shifts occur gradually over time. See Tab. 1 for a comparison. More formally, let $\mathcal{T}=\{1,\dots,T\}$ be a set of $T$ time indices. At test time, let the data at time $t\in\mathcal{T}$ be sampled from a distribution $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})=\mathbb{Q}_{t}({\mathbf{x}})\ \mathbb{Q}_{t}({\textnormal{y}}|{\mathbf{x}})$. The test distributions differ from the source distribution, $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{P}({\mathbf{x}},{\textnormal{y}})\ \forall t>0$, and are non-stationary, meaning $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{Q}_{t^{\prime}}({\mathbf{x}},{\textnormal{y}})$ for $t\neq t^{\prime}$. However, the test distribution evolves gradually such that the discrepancy between consecutive distributions is bounded, $0\leq d(\mathbb{Q}_{t},\mathbb{Q}_{t+1})\leq\epsilon$, where $d$ is a divergence function (Qin et al., 2022). Like in standard TTA, we of course do not observe labels at test time, and hence we observe only a batch of features ${\mathbf{X}}_{t}=\{{\mathbf{x}}_{1,t},\ldots,{\mathbf{x}}_{N,t}\}$, where ${\mathbf{x}}_{n,t}\sim\mathbb{Q}_{t}({\mathbf{x}})$ (i.i.d.). Given the $t$-th batch of features ${\mathbf{X}}_{t}$, the goal is to adapt $f_{\theta}$, forming a new set of parameters $\theta_{t}$ such that $f_{\theta_{t}}$ has better predictive performance on ${\mathbf{X}}_{t}$ than $f_{\theta}$ would have. Since we can only observe features, we assume that the distribution shift must at least take the form of covariate shift: $\mathbb{Q}_{t}({\mathbf{x}})\neq\mathbb{P}({\mathbf{x}})\ \forall t>0$. In addition, a label shift may occur, which poses an additional challenge: $\mathbb{Q}_{t}({\mathbf{y}})\neq\mathbb{P}({\mathbf{y}})\ \forall t>0$.
3 Tracking the Dynamics of Temporal Distribution Shifts
@@
We now present our method: the core idea is that adaptation to temporal distribution shifts can be done by tracking its gradual change in the model’s representations. We employ linear state-space models (SSMs) to capture how test points evolve and drift. The SSM’s cluster representations then serve as an adaptive classification head that evolves with the non-stationarity of the distribution shift. Fig. 1 illustrates our method. In Sec. 3.2, we first introduce the general model and then, in Sec. 3.3, we propose an efficient implementation that leverages the von-Mises-Fisher distribution to model hyperspherical features.
@@ 3.1 Adaptation in Representation Space
Following previous work (Iwasawa & Matsuo, 2021; Boudiaf et al., 2022), we adapt only the last layer of the source model. This lightweight approach is reasonable for Temp-TTA since the distribution shifts only gradually, hence constrained adaptation is needed. From a practical perspective, this circumvents backpropagation through potentially large networks such as foundation models and allows adaptation when only embeddings are provided e.g. by an API. More formally, let the classifier $f_{\theta}$ be a neural network with $L$ total layers. We will treat the first $L-1$ layers, denoted as $f_{\theta}^{L-1}$, as a black box that transforms the original feature vector ${\mathbf{x}}$ into a new (lower-dimensional) representation, which we denote as ${\mathbf{h}}$. The original classifier then maps these representations to the classes as: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{0}{\mathbf{h}}\right)$, where $\text{{softmax}}_{y}\left(\cdot\right)$ denotes the dimension of the softmax’s output corresponding to the $y$-th label index and ${\mathbf{W}}_{0}$ are the last-layer weights. As ${\mathbf{W}}_{0}$ will only be valid for representations that are similar to the training data, we will discard these parameters when performing TempTTA, learning new parameters ${\mathbf{W}}_{t}$ for the $t$-th time step. These new parameters will be used to generate the adapted predictions through the same link function: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{t}{\mathbf{h}}\right)$. In the setting of TempTTA, we observe a batch of features ${\mathbf{X}}_{t}$. Passing them through the model yields corresponding representations ${\mathbf{H}}_{t}$, and this will be the ‘data’ used for the probabilistic model we will describe below. Specifically, we will model how the representations change from ${\mathbf{H}}_{t}$ to ${\mathbf{H}}_{t+1}$ next.
@@ 3.2 A Probabilistic Model of Shift Dynamics
Figure 1: STAD illustrated: It adapts to distribution shifts by inferring dynamic class prototypes ${\mathbf{w}}_{t,k}$ for each class $k$ (different colors) at each test time point. It operates on the representation space of the penultimate layer.
We now describe our general method for a time-evolving adaptive classification head. We assume that, while the representations ${\mathbf{H}}_{t}$ are changing gradually over time, they are still maintaining some class structure in the form of clusters. Our model will seek to track this structure as it evolves. For the intuition of the approach, see Fig. 1. The blue red, and green clusters represent classes of a classification problem. As the distribution shifts from time step $t=1$ to $t=3$, the class clusters shift in representation space. Using latent variables ${\mathbf{w}}_{t,k}$ for the cluster centers, we will assume each representation is drawn conditioned on $K$ latent vectors: ${\bm{h}}_{t,n}\sim p\left({\mathbf{h}}_{t}|{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K}\right)$, where $K$ is equal to the number of classes in the prediction task. After fitting the unsupervised model, the $K$ latent vectors will be stacked to create ${\mathbf{W}}_{t}$, the last-layer weights of the adapted predictive model (as introduced in Sec. 3.1). We now move on to a technical description.
[…]
@@
Let ${\mathbf{H}}_{t}=({\mathbf{h}}_{t,1},\dots,{\mathbf{h}}_{t,N_{t}})\in{\mathbb{R}}^{D\times N_{t}}$ denote the neural representations for $N_{t}$ data points at test time $t$. Let ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})\in{\mathbb{R}}^{D\times K}$ denote the $K$ weight vectors at test time $t$. As discussed above, the weight vector ${\mathbf{w}}_{t,k}$ can be thought of as a latent prototype for class $k$ at time $t$. We denote with ${\mathbf{C}}_{t}=({\mathbf{c}}_{t,1},\dots,{\mathbf{c}}_{t,N_{t}})\in\{0,1\}^{K\times{N_{t}}}$ the $N_{t}$ one-hot encoded latent class assignment vectors ${\mathbf{c}}_{t,n}\in\{0,1\}^{K}$ at time $t$. The $k$-th position of ${\mathbf{c}}_{t,n}$ is denoted with ${\textnormal{c}}_{t,n,k}$ and is $1$ if ${\mathbf{h}}_{t,n}$ belongs to class $k$ and 00 otherwise. Like in standard (static) mixture models, the prior of the latent class assignments $p({\mathbf{c}}_{t,n})$ is a categorical distribution, $p({\mathbf{c}}_{t,n})=\text{{Cat}}(\bm{\pi}_{t})$ with $\bm{\pi}_{t}=(\pi_{t,1},\dots,\pi_{t,K})\in[0,1]^{K}$ and $\sum_{k=1}^{K}\pi_{t,k}=1$. The mixing coefficient $\pi_{t,k}$ gives the a priori probability of class $k$ at time $t$ and can be interpreted as the class proportions. Next, we formally describe how we model the temporal drift of class prototypes.
@@
Fig. 2 depicts the plate diagram of our method. We model the evolution of the $K$ prototypes ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})$ with $K$ independent Markov processes. The resulting transition model is then:
[…]
where $\psi^{\mathrm{trans}}$ denote the parameters of the transition density, notably the transition noise. At each time step, the feature vectors ${\mathbf{H}}_{t}$ are generated by a mixture distribution over the $K$ classes,
[…]
@@
$p($ ${\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})=p({\mathbf{W}}_{1})\prod_{t=1}^{T}p({\mathbf{C}}_{t})p({\mathbf{H}}_{t}|{\mathbf{W}}_{t},{\mathbf{C}}_{t},\psi^{\mathrm{ems}})\prod_{t=2}^{T}p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1},\psi^{\mathrm{trans}})$ (3)
$=\prod_{k}^{K}p({\mathbf{w}}_{1,k})\prod_{t=1}^{T}\prod_{n=1}^{N_{t}}p({\mathbf{c}}_{t,n})\prod_{k=1}^{K}p({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\psi^{\mathrm{ems}})^{{\textnormal{c}}_{t,n,k}}\prod_{t=2}^{T}\prod_{k=1}^{K}p({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\psi^{\mathrm{trans}}).$ (4)
We use the notation ${\mathbf{H}}_{1:T}=\{{\mathbf{H}}_{t}\}_{t=1}^{T}$ to denote the representation vectors ${\mathbf{H}}_{t}$ for all time steps $T$ and analogously for ${\mathbf{W}}_{1:T}$ and ${\mathbf{C}}_{1:T}$. We next outline how we infer the latent class prototypes ${\mathbf{W}}_{1:T}$.
@@
The primary goal is to update the class prototypes ${\mathbf{W}}_{t}$ with the information obtained by the $N_{t}$ representations of test time $t$. At each test time $t$, we are thus interested in the posterior distribution of the prototypes $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:t})$. Once $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:t})$ is known, we can update the classification weights with the new posterior mean. The class weights ${\mathbf{W}}_{t}$ and class assignments ${\mathbf{C}}_{t}$ can be inferred using the Expectation-Maximization (EM) algorithm. In the E-step, we compute $p({\mathbf{W}}_{1:T}{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$. In the M-Step, we maximize the expected complete-data log likelihood with respect to the model parameters:

### v3

@@ Abstract
Distribution shifts between training and test data are inevitable over the lifecycle of a deployed model, leading to performance decay. Adapting a model on test samples can help mitigate this drop in performance. However, most test-time adaptation methods have focused on synthetic corruption shifts, leaving a variety of distribution shifts underexplored. In this paper, we focus on distribution shifts that evolve gradually over time, which are common in the wild but challenging for existing methods, as we show. To address this, we propose STAD, a Bayesian filtering method that adapts a deployed model to temporal distribution shifts by learning the time-varying dynamics in the last set of hidden features. Without requiring labels, our model infers time-evolving class prototypes that act as a dynamic classification head. Through experiments on real-world temporal distribution shifts, we show that our method excels in handling small batch sizes and label shift.
@@ 1 Introduction
Synthetically corrupted images (e.g. CIFAR-10-C) are by far the most commonly used benchmark for assessing progress on TTA, despite concerns about benchmark diversity (Zhao et al., 2023b). These shifts increase the degree of information loss over time, and well-performing TTA methods must learn to preserve a static underlying signal. In this work, we focus on an omnipresent distribution shift of quite a different nature: Temporal distribution shifts encode structural change, not just information loss. Gradual structural change over time is relevant for any deployed model that is operating continuously. While related subfields like gradual domain adaptation (GDA) (Kumar et al., 2020) and temporal domain generalization (TDG) (Bai et al., 2023) are dedicated to these shifts, they have received little attention in TTA. As we will demonstrate using datasets like the Functional Map of the World (FMoW), which classifies land use over time (e.g., rural to urban development), the setting of temporal test-time adaptation (TempTTA) presents significant challenges for existing TTA methods.
To address this gap, we propose State-space Test-time Adaptation (STAD), a method that builds on the power of probabilistic state-space models (SSMs) to represent non-stationary data distributions over time. STAD dynamically adapts a model’s final layer to accommodate an evolving test distribution. Specifically, we employ a probabilistic SSM based on Bayesian filtering to model the evolution of the weight vectors in the final layer, where each vector represents a class, as distribution shift occurs. For generating predictions on newly acquired test batches, we use the SSM’s posterior cluster means as the new parameters. STAD leverages Bayesian updating and does not rely upon normalization mechanisms. As a consequence, STAD excels in scenarios where many TTA methods collapse (Niu et al., 2023), such as adapting with very few samples and under label shift. Our contributions are the following:
  • In Sec.˜2, we detail the setting of TempTTA, which aims to cope with shifts that gradually evolve due to variation in the application domain. Despite being ubiquitous in real-world scenarios, these shifts are understudied in the TTA literature and pose significant challenges to established methods, as we demonstrate in Sec.˜5.1.
  • In Sec.˜3, we propose STAD, a novel method for TempTTA. It adapts to temporal distribution shifts by modeling its dynamics in representation space. No previous work has explicitly modeled these dynamics, which we demonstrate is crucial via an ablation study (Sec.˜5.3).
  • In Sec.˜5, we conduct a comprehensive evaluation of STAD and prominent TTA baselines under authentic temporal shifts. Our results show that STAD excels in this setting (Tab.˜2), yet is applicable beyond temporal shifts providing performance gains on some reproduction datasets (Tab.˜3), synthetic corruptions (Tabs.˜4 and 12) and domain adaptation benchmarks (Tabs.˜9 and 8).
@@ 2 Problem Setting
We focus on the traditional setting of multi-class classification, where $\mathcal{X}\subseteq{\mathbb{R}}^{D}$ denotes the input (feature) space and $\mathcal{Y}\subseteq\{1,\dots,K\}$ denotes the label space. Let $\mathbf{x}$} and y be random variables and $\mathbb{P}({\mathbf{x}},{\textnormal{y}})=\mathbb{P}({\mathbf{x}})\ \mathbb{P}({\textnormal{y}}|{\mathbf{x}})$ the unknown source data distribution. We assume ${\bm{x}}\in\mathcal{X}$ and $y\in\mathcal{Y}$ are realisations of $\mathbf{x}$} and y. The goal of classification is to find a mapping $f_{\theta}$, with parameters $\theta$}, from the input space to the label space $f_{\theta}:\mathcal{X}\to\mathcal{Y}$. Fitting the classifier $f_{\theta}$ is usually accomplished by minimizing an appropriate loss function (e.g. log loss). Yet, our method is agnostic to how $f_{\theta}$ is trained and therefore easy to use with, for instance, a pre-trained model downloaded from the web.
@@
We are interested in adapting a model at test-time to a test distribution that evolves with time. More formally, let $\mathcal{T}=\{1,\dots,T\}$ be a set of $T$ time indices. At test time, let the data at time $t\in\mathcal{T}$ be sampled from a distribution $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})=\mathbb{Q}_{t}({\mathbf{x}})\ \mathbb{Q}_{t}({\textnormal{y}}|{\mathbf{x}})$. The test distributions differ from the source distribution, $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{P}({\mathbf{x}},{\textnormal{y}})\ \forall t>0$, and are non-stationary, meaning $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{Q}_{t^{\prime}}({\mathbf{x}},{\textnormal{y}})$ for $t\neq t^{\prime}$. Like in standard TTA, we of course do not observe labels at test time, and hence we observe only a batch of features ${\mathbf{X}}_{t}=\{{\mathbf{x}}_{1,t},\ldots,{\mathbf{x}}_{N,t}\}$, where ${\mathbf{x}}_{n,t}\sim\mathbb{Q}_{t}({\mathbf{x}})$ (i.i.d.). Given the $t$-th batch of features ${\mathbf{X}}_{t}$, the goal is to adapt $f_{\theta}$, forming a new set of parameters $\theta_{t}$ such that $f_{\theta_{t}}$ has better predictive performance on ${\mathbf{X}}_{t}$ than $f_{\theta}$ would have. Since we can only observe features, we assume that the distribution shift must at least take the form of covariate shift: $\mathbb{Q}_{t}({\mathbf{x}})\neq\mathbb{P}({\mathbf{x}})\ \forall t>0$. In addition, a label shift may occur, which poses an additional challenge: $\mathbb{Q}_{t}({\mathbf{y}})\neq\mathbb{P}({\mathbf{y}})\ \forall t>0$. Temporal shifts, as described above, have been the focus of temporal domain generalization (Bai et al., 2023) and gradual domain adaptation (Abnar et al., 2021). However, both paradigms operate during training, whereas TempTTA is applicable at test time. In Tab.˜1, we contrast TempTTA with adjacent fields highlighting subfields that address temporal shifts. Notably, TempTTA can be seen as a special case of continual TTA (CTTA) (Wang et al., 2022b) with the important distinction that the domain index $t$ is inherently temporal. This is opposed to a categorical domain index (e.g. different corruption types) as is mostly studied in CTTA.
Table 1: Comparison of TempTTA with related fields
Adaptation
Available
Test distribution
stage samples non-stationary time-ordered
Domain generalization (DG) train $\mathbb{P}({\mathbf{x}},{\textnormal{y}})$
Temporal DG train $\mathbb{P}_{1}({\mathbf{x}},{\textnormal{y}}),\dots,\mathbb{P}_{T_{S}}({\mathbf{x}},{\textnormal{y}})$
Domain adaptation (DA) train $\mathbb{P}({\mathbf{x}},{\textnormal{y}}),\mathbb{Q}({\mathbf{x}})$
Gradual DA train $\mathbb{P}({\mathbf{x}},{\textnormal{y}}),\mathbb{Q}_{1}({\mathbf{x}}),\dots,\mathbb{Q}_{T}({\mathbf{x}})$
Test-time training train, test $\mathbb{P}({\mathbf{x}},{\textnormal{y}}),\mathbb{Q}({\mathbf{x}})$
Test-time adaptation (TTA) test $\mathbb{Q}({\mathbf{x}})$
Continual TTA test $\mathbb{Q}_{1}({\mathbf{x}}),\dots,\mathbb{Q}_{T}({\mathbf{x}})$
TempTTA test $\mathbb{Q}_{1}({\mathbf{x}}),\dots,\mathbb{Q}_{T}({\mathbf{x}})$
3 Tracking the Dynamics of Temporal Shifts
@@
We now present our method: the core idea is that adaptation to temporal distribution shifts can be done by tracking its gradual change in the model’s representations. We employ linear SSMs to capture how test points evolve and drift. The SSM’s cluster representations then serve as an adaptive classification head that evolves with the non-stationarity of the distribution shift. Fig.˜1 illustrates our method. In Sec.˜3.2, we first introduce the general model and then, in Sec.˜3.3, we propose an efficient implementation that leverages the von Mises-Fisher distribution to model spherical features.
@@ 3.1 Adaptation in Representation Space
Figure 1: STAD adapts to distribution shifts by inferring dynamic class prototypes ${\mathbf{w}}_{t,k}$ for each class $k$ (different colors) at each test time point. It operates on the representation space of the penultimate layer.
Following previous work (Iwasawa & Matsuo, 2021; Boudiaf et al., 2022), we adapt only the last layer of the source model. This lightweight approach is reasonable for Temp-TTA since the distribution shifts gradually over time, hence constrained adaptation is needed. From a practical perspective, this circumvents backpropagation through potentially large networks such as foundation models and allows adaptation when only embeddings are provided e.g. by an API. More formally, let the classifier $f_{\theta}$ be a neural network with $L$ total layers. We will treat the first L1L-1 layers, denoted as $f_{\theta}^{L-1}$, as a black box that transforms the original feature vector $\mathbf{x}$} into a new (lower-dimensional) representation, which we denote as $\mathbf{h}$}. The original classifier then maps these representations to the classes as: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{0}^{\top}{\mathbf{h}}\right)$, where $\text{{softmax}}_{y}\left(\cdot\right)$ denotes the dimension of the softmax’s output corresponding to the $y$-th label index and ${\mathbf{W}}_{0}$ are the last-layer weights. As ${\mathbf{W}}_{0}$ will only be valid for representations that are similar to the training data, we will discard these parameters when performing TempTTA, learning new parameters ${\mathbf{W}}_{t}$ for the $t$-th time step. These new parameters will be used to generate the adapted predictions through the same link function: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{t}^{\top}{\mathbf{h}}\right)$. In the setting of TempTTA, we observe a batch of features ${\mathbf{X}}_{t}$. Passing them through the model yields corresponding representations ${\mathbf{H}}_{t}$, and this will be the ‘data’ used for the probabilistic model we will describe below. Specifically, we will model how the representations change from ${\mathbf{H}}_{t}$ to ${\mathbf{H}}_{t+1}$ next.
@@ 3.2 A Probabilistic Model of Shift Dynamics
We now describe our general method for a time-evolving adaptive classification head. We assume that, while the representations ${\mathbf{H}}_{t}$ are changing gradually over time, they are still maintaining some class structure in the form of clusters. Our model will seek to track this structure as it evolves. For the intuition of the approach, see Fig.˜1. The blue red, and green clusters represent classes of a classification problem. As the distribution shifts from time step $t=1$ to $t=3$, the class clusters shift in representation space. Using latent variables ${\mathbf{w}}_{t,k}$ for the cluster centers, we will assume each representation is drawn conditioned on $K$ latent vectors: ${\bm{h}}_{t,n}\sim p\left({\mathbf{h}}_{t}|{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K}\right)$, where $K$ is equal to the number of classes in the prediction task. After fitting the unsupervised model, the $K$ latent vectors will be stacked to create ${\mathbf{W}}_{t}$, the last-layer weights of the adapted predictive model (as introduced in Sec.˜3.1). We now move on to a technical description.
@@
Let ${\mathbf{H}}_{t}=({\mathbf{h}}_{t,1},\dots,{\mathbf{h}}_{t,N_{t}})\in{\mathbb{R}}^{D\times N_{t}}$ denote the neural representations for $N_{t}$ data points at test time $t$. Let ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})\in{\mathbb{R}}^{D\times K}$ denote the $K$ weight vectors at test time $t$. As discussed above, the weight vector ${\mathbf{w}}_{t,k}$ can be thought of as a latent prototype for class $k$ at time $t$. We denote with ${\mathbf{C}}_{t}=({\mathbf{c}}_{t,1},\dots,{\mathbf{c}}_{t,N_{t}})\in\{0,1\}^{K\times{N_{t}}}$ the $N_{t}$ one-hot encoded latent class assignment vectors ${\mathbf{c}}_{t,n}\in\{0,1\}^{K}$ at time $t$. The $k$-th position of ${\mathbf{c}}_{t,n}$ is denoted with ${\textnormal{c}}_{t,n,k}$ and is $1$ if ${\mathbf{h}}_{t,n}$ belongs to class $k$ and 0 otherwise. Like in standard (static) mixture models, the prior of the latent class assignments $p({\mathbf{c}}_{t,n})$ is a categorical distribution, $p({\mathbf{c}}_{t,n})=\text{{Cat}}(\bm{\pi}_{t})$ with $\bm{\pi}_{t}=(\pi_{t,1},\dots,\pi_{t,K})\in[0,1]^{K}$ and $\sum_{k=1}^{K}\pi_{t,k}=1$. The mixing coefficient $\pi_{t,k}$ gives the a priori probability of class $k$ at time $t$ and can be interpreted as the class proportions. Next, we formally describe how we model the temporal drift of class prototypes.
@@
We model the evolution of the $K$ prototypes ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})$ with $K$ independent Markov processes. The resulting transition model is
$p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1},\psi^{\mathrm{trans}})=\prod_{k=1}^{K}p({\mathbf{w}}_{t,k}|{\mathbf{w}}_{t-1,k},\psi^{\mathrm{trans}}),\\$ (1)
where $\psi^{\mathrm{trans}}$ denote the parameters of the transition density. At each time step, the feature vectors ${\mathbf{H}}_{t}$ are generated by a mixture distribution over the $K$ classes,
$p({\mathbf{H}}_{t}|{\mathbf{W}}_{t},\psi^{\mathrm{ems}})=\prod_{n=1}^{N_{t}}\sum_{k=1}^{K}\pi_{t,k}\cdot p({\mathbf{h}}_{t,n}|{\mathbf{w}}_{t,k},\psi^{\mathrm{ems}}).$ (2)
@@
$p({\mathbf{H}}_{1:T},{\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})=p({\mathbf{W}}_{1})\prod_{t=2}^{T}p({\mathbf{W}}_{t}|{\mathbf{W}}_{t-1},\psi^{\mathrm{trans}})\prod_{t=1}^{T}p({\mathbf{C}}_{t})p({\mathbf{H}}_{t}|{\mathbf{W}}_{t},{\mathbf{C}}_{t},\psi^{\mathrm{ems}}).$ (3)
We use the notation ${\mathbf{H}}_{1:T}=\{{\mathbf{H}}_{t}\}_{t=1}^{T}$ to denote the representation vectors ${\mathbf{H}}_{t}$ for all time steps $T$ and analogously for ${\mathbf{W}}_{1:T}$ and ${\mathbf{C}}_{1:T}$. A plate diagram of the probabilistic model is depicted in App.˜C. We next outline how we infer the latent class prototypes ${\mathbf{W}}_{1:T}$.
@@
The primary goal is to update the class prototypes ${\mathbf{W}}_{t}$ with the information obtained by the $N_{t}$ representations of test time $t$. At each test time $t$, we are thus interested in the posterior distribution of the prototypes $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:t})$. Once $p({\mathbf{W}}_{t}|{\mathbf{H}}_{1:t})$ is known, we can update the classification weights with the new posterior mean. We can fit the probabilistic model and infer the posterior distribution for the class weights ${\mathbf{W}}_{t}$ and class assignments ${\mathbf{C}}_{t}$ with the Expectation-Maximization (EM) algorithm. In the E-step, we compute the posterior $p({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$. In the M-step, we compute the expectation of the complete-data log-likelihood (Eqn.˜3) with respect to this posterior and then maximize the resulting expression with respect to the model parameters $\phi$: